OpenStudy (anonymous):

Given:\[y''+\frac{2}{r}y'+(1-\frac{l(l+1)}{r^2})y=0\]\[J(r)=r^{\frac{1}{2}}y(r)\] Show that\[J''+\frac{1}{r}J'+(1-\frac{l(l+1)+\frac{1}{4}}{r^2})J=0\] I can get to:\[J''+\frac{1}{r}J'=r^\frac{1}{2}((\frac{l(l+1)}{r^2}-1)y+\frac{1}{4r^2}y)\]

OpenStudy (anonymous):

Ok, it looks like your method has overcomplicated things, so I will use a different method. To begin, write \(y=Jr^{-\frac{1}{2}}\) and calculate the derivatives of y in terms of J: \[\large y=Jr^{-\frac{1}{2}}\] \[\large y'=J'r^{-\frac{1}{2}}-\frac{1}{2}Jr^{-\frac{3}{2}}\] \[\large y''=J''r^{-\frac{1}{2}}-J'r^{-\frac{3}{2}}+\frac{3}{4}Jr^{-\frac{5}{2}}\]And then substitute these into your first equation.

OpenStudy (anonymous):

Do you understand the differentiation? Are you comfortable with this method; do you not need explanation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

To further simplify calculations, I would write \(1-\frac{l(l+1)}{r^2}=q\).

OpenStudy (anonymous):

You should check my calculations yourself. The rest is pretty simple. Can you figure it out on your own?

OpenStudy (anonymous):

Yes, give me a minute, I'll go through everything this way. I'll post results

OpenStudy (anonymous):

This worked alot better than what I was doing. I got:\[J''+\frac{1}{r}J'+(1+\frac{\frac{1}{4}+l(l+1)}{r^2})J=0\]So I just made a sign mistake somewhere.